Greta, an elderly investor, has a degree of risk aversion of A = 4 when applied to return on wealth over a one-year horizon. She is pondering two portfolios, the S&P/TSX Composite Index and a hedge fund, as well as a number of one-year strategies. (All rates are annual and continuously compounded.) The S&P/TSX Composite risk premium is estimated at 7% per year, with a SD of 18%. The hedge fund risk premium is estimated at 5% with a SD of 25%. The returns on both of these portfolios in any particular year are uncorrelated with its own returns in other years. They are also uncorrelated with the returns of the other portfolio in other years. The hedge fund claims the correlation coefficient between the annual return on the S&P/TSX Composite and the hedge fund return in the same year is zero, but Greta is not fully convinced by this claim.


What should be Greta’s capital allocation? Assume that the correlation between the annual returns on the two portfolios is indeed zero. (Do not round intermediate calculations. Enter your answers rounded to 2 decimal places.)

S&P %
Hedge %
Risk-free asset %
Answer #1

Y = Risk Premium Aaa S&P Y = - = 0.07 4x (0.18) 0.54 / 54% Hedge - Y = 0.05 4X(0.252 = 0.20 / 20% Risk Free Asset = -0.54 - 0capital has been allocated in such a way that it produces maximum utility to the investor given her risk aversion.